​Claim: Fewer than 91​% of adults have a cell phone. In a reputable poll of 1079 ​adults, 83​% said that they have a cell phone. Find the value of the test statistic.
Answer #1

Solution :

The null and alternative hypothesis is

H0 : p = 0.91

Ha : p < 0.91

n = 1079

= 0.83

P0 = 0.91

1 - P0 = 1 - 0.91 = 0.09

Test statistic = z =

= - P0 / [P0 * (1 - P0 ) / n]

= 0.83 - 0.91 / [(0.91 * 0.09) / 1079 ]

Test statistic = z = -9.18